How many milliliters of 11.2 M nitric acid solution should be used to prepare 5.00 L of 0.500 M HNO3?
1
Expert's answer
2020-04-18T07:07:52-0400

C1=11.2 M

V2=5.00 L

C2=0.5 M

V1 -?


C1*V1=C2*V2;

V1=C2*V2/C1=0.5*5/11.2=0.22321 L=223.21 mL;

Answer: 0.22321 L or 223.21 mL of 11.2 M nitric acid solution should be used